# Logistic-Beta Processes for Dependent Random Probabilities with Beta Marginals

### v2

@@ 5.2 Application to pregnancy outcome data analysis
We use the logistic-beta dependent Dirichlet process mixture model (12)–(13) to analyze DDE exposure and pregnancy outcome data (Longnecker et al.,, 2001). DDE is a metabolite of chemical compound DDT that was widely used as an insecticide. We analyze how the conditional distribution of gestational age of delivery $y$ changes with DDE exposure level $x$, focusing on the probability of preterm birth, defined as gestational age of delivery less than 37 weeks. The dataset is publicly available in R package BNPmix (Corradin et al.,, 2021). For illustration we focus on the $n=1023$ women who smoked during pregnancy; results for other groups are in Appendix D.2.
@@
Since there is no publicly available software for a dependent Dirichlet process with covariate-dependent weights, we compare with the logit stick-breaking process (Ren et al.,, 2011). We choose basis functions as in Rigon and Durante, (2021) by considering normalized feature map kernels obtained from a natural cubic spline basis with 6 degrees of freedom. In both our model and logit stick-breaking, we use a mixture of linear regressions with covariate-dependent weights as in (12). We compare the results in terms of the probability of preterm birth given DDE exposure. For the logistic-beta dependent Dirichlet process, we consider three different settings of parameter $b\in\{1/5,1,2\}$, referred to as settings 1, 2, and 3, respectively. This corresponds to prior co-clustering probabilities for any $x$ of $5/6$, $1/2$, and $1/3$. For the logit stick-breaking process, we choose hyperparameters such that prior co-clustering probabilities closely match with the proposed model within each setting.
@@
Posterior computation is carried out with Gibbs samplers for both models. We obtained 35,000 iterations, with the first 5,000 samples discarded as burn-in. Both algorithms were implemented in R and had similar running times; 7.5 minutes for the logistic-beta dependent Dirichlet process model and 5.6 minutes for the logit stick-breaking process model under setting 2.
@@
Figure 5: Estimated probability of preterm birth with 95% credible intervals against DDE exposure level from the logistic-beta dependent Dirichlet process model (top) and the logit stick-breaking process model (bottom) with three different hyperparameter settings. DDE exposure data are represented by small vertical bars.
@@
The estimated preterm birth probabilities given DDE exposures are summarized in Fig. 5. For both models, the probabilities tend to increase with DDE levels, with wider credible intervals at high levels due to the increasingly sparse data. There is a stark difference between the two models in terms of how the estimates and credible intervals vary according to hyperparameter choice. Although each hyperparameter setting yields similar prior properties and both models use the same basis functions to incorporate dependence in $x$, Fig. 5 shows that the posterior estimate of the logit stick-breaking process is highly sensitive compared to the proposed model. Specifically, the hyperparameter choice of the logit stick-breaking process model has a substantial impact on uncertainty estimates at high DDE exposure levels, where setting 1 and setting 3 of the logit stick-breaking process model have much wider and narrower credible intervals, respectively, compared to the proposed model. This sensitivity, as noted in Wade et al., (2023), makes conclusions drawn from the logit stick-breaking process model questionable, whereas the proposed logistic-beta dependent Dirichlet process model is much more robust to the choice of hyperparameters.
@@
Appendix D.2 contains details on pregnancy outcome data analysis settings and additional results with the non-smoking group of size $n=1290$ that gives the same conclusion.
@@ 6 Discussion
Although we have focused on two particular applications, logistic-beta processes can be usefully integrated within many other Bayesian nonparametric models. A particularly interesting case is to incorporate covariate dependence into Indian buffet processes, relying on a beta-Bernoulli representation. An interesting application is to species distribution modeling in ecology (Elith and Leathwick,, 2009), where the features correspond to species and covariates correspond to environmental data, such as location or temperature. Such a model can accommodate and predict new species discovery, which is common in ecological studies of insect and fungi.
@@ Acknowledgements
The research of Changwoo Lee and Huiyan Sang was partially supported by the U.S. National Science Foundation (2210456 and 2220231). The research of Alessandro Zito and David Dunson was supported by U.S. Office of Naval Research (N00014-21-1-2510-P00004-7), and National Institutes of Health Grant (R01ES035625). This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (856506).
Supplementary Material
@@
Appendix A Details of adaptive Pólya proposal
We describe the adaptive Pólya proposal scheme for sampling $\lambda$ from a distribution that involves Pólya prior density ${\pi_{\textsc{po}}}(\lambda;a,b)$, such as expression (9) in Step 2 of Algorithm 9. The evaluation of Pólya density ${\pi_{\textsc{po}}}$ is problematic since it is expressed in terms of alternating series whose evaluation is numerically unstable; see equation (3). In the context of regularized logistic regression where the same problem arises, Gramacy and Polson, (2012) proposed using independent Metropolis-Hastings with Pólya prior proposal $\mathrm{Polya}(a,b)$ to avoid the evaluation of ${\pi_{\textsc{po}}}$ as these cancel out in the acceptance ratio. However, such a sampler ignores all information in the data and can suffer from a low acceptance rate.
The adaptive Pólya proposal scheme selects the proposal distribution as $\mathrm{Polya}(a^{\prime},b^{\prime})$ with suitably chosen $(a^{\prime},b^{\prime})$ pair such that $a^{\prime}+b^{\prime}=a+b=c$, leveraging the identity in Proposition 1 to avoid evaluation of the Pólya density ${\pi_{\textsc{po}}}$. To see this, for example with Metropolis-Hastings, a new candidate is drawn from $\lambda^{\star}\sim\mathrm{Polya}(a^{\prime},b^{\prime})$, and accepted with probability $\min\{1,\exp\{(\lambda-\lambda^{\star})(ab-a^{\prime}b^{\prime})/2\}g(\lambda^{\star})/g(\lambda)\}$ where ${\pi_{\textsc{po}}}$ terms are cancelled out; here $g(\lambda)$ corresponds to multivariate normal density in (9) as a function of $\lambda$. This adaptive Pólya proposal scheme can be used in more complex transition kernels, such as a particle Gibbs sampler (Andrieu et al.,, 2010), see Appendix C.2 for examples.
There are many ways to incorporate previous samples’ information in the choice of $(a^{\prime},b^{\prime})$. An example is the moment matching method with a running average of $\lambda$, denoted as $\bar{\lambda}$, that is recursively updated as $\bar{\lambda}\leftarrow\bar{\lambda}(m-1)/m+\lambda/m$ in $m$th iteration of the algorithm, leading to adaptive Metropolis-within-Gibbs (Roberts and Rosenthal,, 2009). Assuming $a^{\prime}\leq b^{\prime}$ without loss of generality due to the symmetry of the Pólya distribution in its arguments, which implies $a^{\prime}\leq c/2$, the $\mathrm{Polya}(a^{\prime},c-a^{\prime})$ has a mean $h(a^{\prime})=2\{\psi(a^{\prime})-\psi(c-a^{\prime})\}/(2a^{\prime}-c)$ (or $h(a^{\prime})=2\psi^{\prime}(c/2)$ if $a^{\prime}=c/2$) that is decreasing on $a^{\prime}\in(0,c/2]$. Thus, we set $a^{\prime}=(a+b)/2$ if $\bar{\lambda}$ is less than or equal to $2\psi^{\prime}((a+b)/2)$, and otherwise we set $a^{\prime}$ to the solution of $h(x)=\bar{\lambda}$ which can be easily obtained with numerical optimization algorithms.
Appendix B Proofs and additional propositions
B.1 Properties of Pólya distributions
@@
We collect results related to the Pólya distribution, including its moments (Proposition B.1) and density identity (Proposition 1).
Proposition B.1.
@@
Next, we present a proof of Proposition 1, an identity between Pólya densities.
@@
$=\frac{B(a,b)}{B(a^{\prime},b^{\prime})}\sum_{k=0}^{\infty}(-1)^{k}\binom{a+b+k-1}{k}\frac{(k+(a+b)/2)}{B(a,b)}\exp\left(-\frac{(k+a)(k+b)}{2}\lambda+\frac{(ab-a^{\prime}b^{\prime})\lambda}{2}\right)$
@@
which proves Proposition 1. ∎
B.2 Properties of multivariate logistic-beta and logistic-beta processes
We summarize results related to multivariate logistic-beta and logistic-beta processes, including the moments (Proposition 2), correlation range (Corollary 1), and representations of logistic-beta processes under normalized feature map kernel (Proposition 3).
Below we present proofs of Proposition 2 and Corollary 1.
@@ B.2 Properties of multivariate logistic-beta and logistic-beta processes
which shows Proposition 2. Considering a bivariate case, if $a=b$, the range of $\operatorname*{corr}(\eta_{1},\eta_{2})$ is $[-1,1]$. If $a\neq b$, the upper bound 1 of $\operatorname*{corr}(\eta_{1},\eta_{2})$ is attained at $R_{12}=1$, and lower bound of $1-4\{\psi(a)-\psi(b)\}/\{(a-b)(\psi^{\prime}(a)+\psi^{\prime}(b))\}$ is attained at $R_{12}=-1$. This shows Corollary 1. ∎
Next, we present a proof of Proposition 3.
@@
Consider a normalized feature map $\phi:\mathscr{X}\to\mathds{R}^{q}$ and the corresponding correlation kernel $\mathcal{R}(x,x^{\prime})=\langle\phi(x),\phi(x^{\prime})\rangle$ for logistic-beta process. Let $\Phi=[\phi_{k}(x_{i})]_{ik}$ be an $n\times q$ basis matrix with $i$th row corresponding to $\phi(x_{i})$. Then, for a realization $\eta=(\eta(x_{1}),\dots,\eta(x_{n}))^{\mathrm{\scriptscriptstyle T}}$, we have
$\eta\mid\lambda\sim N_{n}\left(0.5\lambda(a-b)1_{n},\lambda\Phi\Phi^{\mathrm{\scriptscriptstyle T}}\right),\quad\lambda\sim\mathrm{Polya}(a,b).$
However, $\eta\mid\lambda$ is a rank deficient multivariate normal when $q<n$. By introducing $\gamma\sim N_{q}(0_{q},I_{q})$, it can be equivalently written as
$\eta=0.5\lambda(a-b)1_{n}+\lambda^{1/2}\Phi\gamma,\quad\lambda\sim\mathrm{Polya}(a,b),\quad\gamma\sim N_{q}(0_{q},I_{q}),$
which corresponds to the hierarchical representation of Proposition 3. Also,
$\eta$ $=0.5\lambda(a-b)1_{n}+\lambda^{1/2}\Phi\gamma$
$=0.5\lambda(a-b)1_{n}-0.5\lambda(a-b)\Phi 1_{q}+0.5\lambda(a-b)\Phi 1_{q}+\lambda^{1/2}\Phi\gamma$
$=0.5\lambda(a-b)(1_{n}-\Phi 1_{q})+\Phi\{0.5\lambda(a-b)1_{q}+\lambda^{1/2}\gamma\}.$
@@
$\eta=\{\psi(a)-\psi(b)\}(1_{n}-\Phi 1_{q})+\Phi\beta,\quad\beta\sim{\textsc{lb}}(a,b,I_{q}),$
where we used $E(\lambda)=2\{\psi(a)-\psi(b)\}/(a-b)$ as in Proposition B.1. This shows the linear predictor representation of Proposition 3. ∎
B.3 Logistic-beta dependent Dirichlet process and related models
We present results on the logistic-beta dependent Dirichlet process, including full weak support property (Theorem B.1) and correlation between random probability measures (Theorem 1). We first present proof of Theorem 1, which analyzes prior correlation structure induced by logistic-beta dependent Dirichlet process.
@@ B.3 Logistic-beta dependent Dirichlet process and related models
The proof closely follows that for Theorem 1 of Dunson and Park, (2008). Recall that we are considering a single-atoms logistic-beta dependent Dirichlet process, where the atom process does not depend on $x$. Denote the base measure $G_{0}$ and $V_{h}(x)=\sigma\{\eta_{h}(x)\}$, where marginally $V_{h}(x)\sim\mathrm{Beta}(1,b)$. Since the dependent Dirichlet process prior implies $G_{x}$ marginally follows a Dirichlet process with concentration parameter $b$, we have $E\{G_{x}(B)\}=G_{0}(B)$ and $\operatorname*{var}\{G_{x}(B)\}=G_{0}(B)\{1-G_{0}(B)\}/(1+b)$ (Müller et al.,, 2015). In addition to the notation $\mu(x_{i},x_{j})=E[\sigma\{\eta(x_{i})\}\sigma\{\eta(x_{j})\}]$ where $\eta(\cdot)\sim{\textsc{lbp}}(1,b,\mathcal{R})$, denote $\mu(x)=E[\sigma\{\eta(x)\}]$ which is simply $\mu(x)=1/(1+b)$ for any $x$. The mixed moment is
$E\{G_{x_{i}}(B)G_{x_{j}}(B)\}$
=\,\,= $E\left[\sum_{h=1}^{\infty}V_{h}(x_{i})\left\{\prod_{l=1}^{h-1}(1-V_{l}(x_{i}))\right\}\delta_{\theta_{h}}(B)\times\sum_{h=1}^{\infty}V_{h}(x_{j})\left\{\prod_{l=1}^{h-1}(1-V_{l}(x_{j}))\right\}\delta_{\theta_{h}}(B)\right]$
=\,\,= $E\left[\sum_{h=1}^{\infty}V_{h}(x_{i})V_{h}(x_{j})\prod_{l=1}^{h-1}(1-V_{h}(x_{i}))(1-V_{h}(x_{j}))\delta_{\theta_{h}}(B)\right]+$
$E\left[\sum_{h=1}^{\infty}\sum_{l=1}^{h-1}V_{h}(x_{i})\prod_{r=1}^{l-1}\left\{(1-V_{r}(x_{i}))(1-V_{r}(x_{j}))\right\}\{V_{l}(x_{j})-V_{l}(x_{i})V_{l}(x_{j})\}\right.$
×$\left.\times\prod_{s={l+1}}^{h-1}\{1-V_{s}(x_{i})\}\delta_{\theta_{h}}(B)\delta_{\theta_{l}}(B)\right]+$
$E\left[\sum_{h=1}^{\infty}\sum_{l=h+1}^{\infty}V_{l}(x_{j})\prod_{r=1}^{h-1}\left\{(1-V_{r}(x_{i}))(1-V_{r}(x_{j}))\right\}\{V_{h}(x_{i})-V_{h}(x_{i})V_{h}(x_{j})\}\right.$
×$\left.\times\prod_{s={h+1}}^{l-1}\{1-V_{s}(x_{j})\}\delta_{\theta_{h}}(B)\delta_{\theta_{l}}(B)\right]$
=\,\,= $\sum_{h=1}^{\infty}\mu(x_{i},x_{j})\{1-\mu(x_{i})-\mu(x_{j})+\mu(x_{i},x_{j})\}^{h-1}G_{0}(B)+$
$(\mu(x_{j})-\mu(x_{i},x_{j}))\sum_{l=1}^{\infty}\{1-\mu(x_{i})-\mu(x_{j})+\mu(x_{i},x_{j})\}^{l-1}G_{0}(B)^{2}+$
$(\mu(x_{i})-\mu(x_{i},x_{j}))\sum_{h=1}^{\infty}\{1-\mu(x_{i})-\mu(x_{j})+\mu(x_{i},x_{j})\}^{h-1}G_{0}(B)^{2}$
=\,\,= $\frac{\mu(x_{i},x_{j})(G_{0}(B)-G_{0}(B)^{2})}{\mu(x_{i})+\mu(x_{j})-\mu(x_{i},x_{j})}-G_{0}(B)^{2}$
which implies, using $\mu(x)=1/(1+b)$ for any $x$,
$\operatorname*{corr}\{G_{x_{i}}(B),G_{x_{j}}(B)\}$ $=\frac{\mu(x_{i},x_{j})(G_{0}(B)-G_{0}(B)^{2})}{(\mu(x_{i})+\mu(x_{j})-\mu(x_{i},x_{j}))\times(G_{0}(B)(1-G_{0}(B))/(1+b))}$
$=\frac{(1+b)\mu(x_{i},x_{j})}{2/(1+b)-\mu(x_{i},x_{j})}=\frac{(1+b)^{2}}{2/\mu(x_{i},x_{j})-(1+b)}$
and this result not only holds for single-atoms logistic-beta dependent Dirichlet process, but also for general single-atoms dependent Dirichlet processes with $\mu(x_{i},x_{j})=E\{V(x_{i})V(x_{j})\}$ where $V(x)$ is a distribution of stick-breaking ratios so that $V_{h}(x)$ are independent copies across all $h\in\mathds{N}$. ∎
Next, we present the prior correlation structure under other various dependent Dirichlet process models. In addition to the logistic-beta dependent Dirichlet process (M1), we provide a derivation of correlation lower bounds under three different dependent Dirichlet process models: (M2) squared AR(1) process-based dependent Dirichlet process (DeYoreo and Kottas,, 2018), (M3) time-series dependent Dirichlet process (Nieto-Barajas et al.,, 2012), and (M4) copula-based dependent Dirichlet process. Similar to the logistic-beta dependent Dirichlet process, these models induce dependence through a stochastic process $V(\cdot)$ on stick-breaking ratios that have beta marginals but with different constructions.
First, from DeYoreo and Kottas, (2018, Appendix A.1.), the quantity $\operatorname*{corr}\{V(x_{i}),V(x_{j})\}$ under the squared ar(1) process-based dependent Dirichlet process has the greatest lower bound of $b^{1/2}(b+1)(b+2)^{1/2}-b(b+2)$ when autoregressive parameter converges to 0. This implies $E\{V(x_{i})V(x_{j})\}$ has the infimum of $\{b^{3/2}+(1-b)(b+2)^{1/2}\}/\{(b+1)(b+2)^{1/2}\}$, using $E\{V(x)\}=1/(b+1)$ and $\operatorname*{var}\{V(x)\}=b/((b+1)^{2}(b+2))$ for any $x$ under beta distributed $V(x)$. This leads to the greatest lower bound of correlation between random probability measures $(1+b)/([2(b+2)^{1/2}/\{b^{3/2}+(1-b)(b+2)^{1/2}\}]-1)$, which converges to $0.75$ as $b\to\infty$.
Next, the infimum of $\operatorname*{corr}\{V(x_{i}),V(x_{j})\}$ under the time-series dependent Dirichlet process (Nieto-Barajas et al.,, 2012) is when $V(x_{i})$ and $V(x_{j})$ are independent so that $\inf\operatorname*{corr}\{V(x_{i}),V(x_{j})\}=0$. Thus, $E\{V(x_{i})V(x_{j})\}$ has the infimum of $1/(b+1)^{2}$, which leads to the infimum of correlation between random probability measures $\inf\operatorname*{corr}\{G_{x_{i}}(B),G_{x_{j}}(B)\}=(1+b)/(1+2b)$. This fact is also noted in Nieto-Barajas et al., (2012) as a special case when stick-breaking ratios are independent, and the same fact also applies to the dependent Dirichlet process proposed by Taddy, (2010) when the autoregressive parameter converges to 0.
Finally, we analyze a copula-based single-atoms dependent Dirichlet process, with beta stick-breaking ratios constructed through a copula that induces a stochastic process with $\mathrm{Beta}(1,b)$ marginals. Finding infimum of $\operatorname*{corr}\{G_{x_{i}}(B),G_{x_{j}}(B)\}$ is equivalent to finding the infimum of correlation between stick-breaking ratios $\operatorname*{corr}\{V(x_{i}),V(x_{j})\}$, which is achieved at the Fréchet lower bound copula (Embrechts et al.,, 2002). In other words, minimum correlation is attained when $V(x_{i})$ and $V(x_{j})$ satisfy $V(x_{j})=F_{B}^{-1}(1-F_{B}(V(x_{i})))$ (counter-monotonic), where $F_{B}$ is the c.d.f. of $\mathrm{Beta}(1,b)$.
Although the analytic form of $\inf\operatorname*{corr}\{V(x_{i}),V(x_{j})\}$ under common atoms logistic-beta dependent Dirichlet process and copula-based dependent Dirichlet process are not readily available due to the nonlinear transformations involved, those can be estimated with Monte Carlo simulation where we used 5,000 simulated datasets of size 1,000 to calculate the correlation estimates.
Figure B.1: Illustration of different correlation lower bounds under four different single-atoms dependent Dirichlet process. (Left) Infimum of correlation between random probability measures under models M1 (solid, black), M2 (dashed, red), M3 (dotted, purple), and M4 (dot-dashed, blue). (Right) Scatterplot of bivariate betas with $\mathrm{Beta}(1,2)$ marginals, corresponding to circles of the left panel.
Figure B.1 illustrates different correlation lower bounds under four different single-atoms dependent Dirichlet processes with concentration parameter $b=2$. The difference in the induced dependence structure is clearly seen from scatterplots, where 1,000 random samples are drawn from a corresponding bivariate beta distribution that induces minimum correlation from models M1–M4. The Fréchet lower bound clearly shows counter-monotonicity between two beta random variables, which induces minimum possible correlation. Transformation of logistic-beta also can induce a negative correlation between beta random variables, while others cannot. In summary, logistic-beta dependent Dirichlet processes can capture a wide range of dependence when $b$ is small to moderate.
The following theorem describes the full-week support property of logistic-beta dependent Dirichlet processes.
Theorem B.1.
Consider a logistic-beta dependent Dirichlet process (10) with an atom process $\{\theta(x):x\in\mathscr{X}\}$ with support $\Theta$ that can be represented with a collection of copulas with positive density w.r.t. Lebesgue measure. Then, the logistic-beta dependent Dirichlet process has full weak support, i.e. the topological support of the logistic-beta dependent Dirichlet process coincides with the space of collections of all probability measures with support $\Theta$ indexed by $x$.

### v3

@@ 5.3 Application to pregnancy outcome data analysis
    We use the logistic-beta dependent Dirichlet process mixture model (14)–(15) to analyze DDE exposure and pregnancy outcome data (Longnecker et al.,, 2001). DDE is a metabolite of the chemical compound DDT that was widely used as an insecticide. We analyze how the conditional distribution of gestational age of delivery $y$ changes with the level of exposure to DDE $x$, focusing on the probability of preterm birth, defined as gestational age of delivery less than 37 weeks. The dataset is publicly available in R package BNPmix (Corradin et al.,, 2021). For illustration, we focus on the $n=1023$ women who smoked during pregnancy; results for other groups are in Section S.3.3 of the Appendix.
@@
    We compare the result with the LSBP mixture model in terms of the estimated probability of preterm birth given DDE exposure $\operatorname*{pr}(y\leq 37\mid x)$ as a function of $x$. For the choice of correlation kernel $\mathcal{R}$ for LB-DDP and basis functions for LSBP, similar to Rigon and Durante, (2021), we consider normalized feature map kernels obtained from a natural cubic spline basis with 6 degrees of freedom. In both models, the atom processes are specified as in (14), leading to a mixture of linear regressions with covariate-dependent weights. For the choice of hyperparameters, we consider the same three settings $b\in\{1/5,1,2\}$ as in the simulation study, denoted setting 1, 2, 3, respectively, where the LSBP hyperparameters are chosen such that the prior tie probabilities closely match the proposed model within each setting.
@@
    Posterior computation is performed with Gibbs samplers for both models. We obtained 35,000 iterations and the first 5,000 samples were discarded as burn-in. Both algorithms were implemented in R and had similar running times; 7.5 minutes for the LB-DDP mixture model and 5.6 minutes for the LSBP mixture model in setting 2.
@@
    Figure 6: Estimated probability of preterm birth with 95% credible intervals against DDE exposure level from the LB-DDP mixture model (top) and the LSBP mixture model (bottom) with three different hyperparameter settings corresponding to prior tie probabilities approximately 0.84, 0.50, and 0.34. DDE data are represented by small vertical bars.
@@
    The estimated preterm birth probabilities given DDE exposures are summarized in Figure 6. For both models, the probabilities tend to increase with DDE levels, with wider credible intervals at high levels because of the increasingly sparse data. There is a stark difference between the two models in terms of how the estimates and credible intervals vary according to hyperparameter choice. Although each hyperparameter setting yields similar prior properties and both models use the same basis functions to incorporate dependence in $x$, Figure 6 shows that the posterior estimate of the LSBP is highly sensitive compared to the proposed model. Specifically, the hyperparameter choice of the LSBP model has a substantial impact on uncertainty estimates at high DDE exposure levels, where setting 1 and setting 3 of the LSBP model have much wider and narrower credible intervals, respectively, compared to the proposed model. This sensitivity, as noted in Wade and Inacio, (2025), makes conclusions drawn from the LSBP model questionable, whereas the proposed LB-DDP model is much more robust to the choice of hyperparameters.
@@
    Section S.3.3 of the Appendix contains details on pregnancy outcome data analysis settings and additional results with the non-smoking group of size $n=1290$ that gives the same conclusion.
@@ 6 Discussion
    Although we have focused on two particular applications, logistic-beta processes can be usefully integrated within many other Bayesian nonparametric models. A particularly intriguing case is to incorporate covariate dependence into Indian buffet processes, relying on a beta-Bernoulli representation. An interesting application is to species distribution modeling in ecology (Elith and Leathwick,, 2009), where the features correspond to species and covariates correspond to environmental data, such as location or temperature. Such a model can accommodate and predict new species discovery, which is common in ecological studies of insect and fungi.
    One promising future avenue of research is to develop a new family of stochastic processes having possibly different beta marginal distributions, while maintaining computational benefits when used for modeling dependent random probabilities. In the logistic-beta process, the common beta marginal property is due to the shared Pólya mixing parameter $\lambda$. Such a new family could be used to develop even more flexible dependent Bayesian nonparametric models, such as DDP models with a covariate-dependent concentration parameter.
    Software
@@ Acknowledgements
    This project has received funding from U.S. National Science Foundation (DMS-2210456, DMS-2220231, IIS-2426762), U.S. National Institutes of Health (R01ES035625), U.S. Office of Naval Research (N00014-21-1-2510-P00004-7), and from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (856506).
@@
    Appendix S.1 Proofs and additional propositions
    S.1.1 Properties of Pólya distributions
@@
    We collect results related to the Pólya distribution, including its moments (Proposition S.1.1) and density identity (Proposition 2.1).
    Proposition S.1.1.
@@
    S.1.1.1 Proof of Proposition 2.1
@@
    $=\frac{B(a,b)}{B(a^{\prime},b^{\prime})}\sum_{k=0}^{\infty}\left\{(-1)^{k}\binom{a+b+k-1}{k}\frac{(k+(a+b)/2)}{B(a,b)}\times\right.$
    $\qquad\qquad\qquad\quad\left.\exp\left(-\frac{(k+a)(k+b)}{2}\lambda+\frac{(ab-a^{\prime}b^{\prime})\lambda}{2}\right)\right\}$
@@
    which proves Proposition 2.1. ∎
    S.1.2 Properties of multivariate logistic-beta and logistic-beta processes
    We summarize results related to multivariate logistic-beta and logistic-beta processes, including the moments (Proposition 2.2), correlation range (Corollary 2.1), and representations of logistic-beta processes under normalized feature map kernel (Proposition 2.3).
    S.1.2.1 Proofs of Proposition 2.2 and Corollary 2.1
@@
    which shows Proposition 2.2. Considering a bivariate case, if $a=b$, the range of $\operatorname*{corr}(\eta_{1},\eta_{2})$ is $[-1,1]$. If $a\neq b$, the upper bound 1 of $\operatorname*{corr}(\eta_{1},\eta_{2})$ is attained at $R_{12}=1$, and lower bound of $1-4\{\psi(a)-\psi(b)\}/\{(a-b)(\psi^{\prime}(a)+\psi^{\prime}(b))\}$ is attained at $R_{12}=-1$. This shows Corollary 2.1. ∎
    S.1.2.2 Proof of Proposition 2.3
@@
    Consider a normalized feature map $\bm{\phi}:\mathscr{X}\to\mathds{R}^{q}$ and the corresponding correlation kernel $\mathcal{R}(x,x^{\prime})=\langle\bm{\phi}(x),\bm{\phi}(x^{\prime})\rangle$ for logistic-beta process. Let $\bm{\Phi}=[\phi_{k}(x_{i})]_{ik}$ be an $n\times q$ basis matrix with $i$th row corresponding to $\bm{\phi}(x_{i})$. Then, for a realization $\bm{\eta}=(\eta(x_{1}),\dots,\eta(x_{n}))^{\mathrm{\scriptscriptstyle T}}$, we have
    $\bm{\eta}\mid\lambda\sim N_{n}\left(0.5\lambda(a-b)\bm{1}_{n},\lambda\bm{\Phi}\bm{\Phi}^{\mathrm{\scriptscriptstyle T}}\right),\quad\lambda\sim\mathrm{Polya}(a,b).$
    However, $\bm{\eta}\mid\lambda$ is a rank deficient multivariate normal when $q<n$. By introducing $\bm{\gamma}\sim N_{q}(\bm{0}_{q},\mathbf{I}_{q})$, it can be equivalently written as
    $\eta=0.5\lambda(a-b)\bm{1}_{n}+\lambda^{1/2}\bm{\Phi}\bm{\gamma},\quad\lambda\sim\mathrm{Polya}(a,b),\quad\bm{\gamma}\sim N_{q}(\bm{0}_{q},\mathbf{I}_{q}),$
    which corresponds to the hierarchical representation of Proposition 2.3. Also,
    $\bm{\eta}$ $=0.5\lambda(a-b)\bm{1}_{n}+\lambda^{1/2}\bm{\Phi}\bm{\gamma}$
    $=0.5\lambda(a-b)\bm{1}_{n}-0.5\lambda(a-b)\bm{\Phi}\bm{1}_{q}+0.5\lambda(a-b)\bm{\Phi}\bm{1}_{q}+\lambda^{1/2}\bm{\Phi}\bm{\gamma}$
    $=0.5\lambda(a-b)(\bm{1}_{n}-\bm{\Phi}\bm{1}_{q})+\bm{\Phi}\{0.5\lambda(a-b)\bm{1}_{q}+\lambda^{1/2}\bm{\gamma}\}.$
@@
    $\bm{\eta}=\{\psi(a)-\psi(b)\}(\bm{1}_{n}-\bm{\Phi}\bm{1}_{q})+\bm{\Phi}\bm{\beta},\quad\bm{\beta}\sim{\textsc{lb}}(a,b,\mathbf{I}_{q}),$
    where we used $E(\lambda)=2\{\psi(a)-\psi(b)\}/(a-b)$ as in Proposition S.1.1. This shows the linear predictor representation of Proposition 2.3. ∎
    S.1.3 Logistic-beta dependent Dirichlet process and related models
    We present results on the logistic-beta dependent Dirichlet process, including full weak support property (Theorem 4.1), correlation between random probability measures (Theorem 4.2), and continuity (Theorem 4.4).
    S.1.3.1 Proof of Theorem 4.1
@@
    First consider the single-atoms logistic-beta dependent Dirichlet process, so that atoms $\theta_{h}$ are i.i.d. samples from the base measure $G_{0}$. Let $\eta(\cdot)\sim{\textsc{lbp}}(1,b,\mathcal{R})$ and $F_{B}$ be the c.d.f. of $\mathrm{Beta}(1,b)$. Consider an $n$-dimensional continuous random variable $\mathbf{U}=(F_{B}[\sigma\{\eta(x_{1})\}],\dots,F_{B}[\sigma\{\eta(x_{n})\}])$ where its cumulative distribution $C_{x_{1},\dots,x_{n}}$ is a copula since $\mathbf{U}$ has uniform marginals by construction. It is clear that $\mathbf{U}$ has a positive density w.r.t. Lebesgue measure since $F_{B}$ and $\sigma$ are all continuous and injective. Now consider a collection of copula functions $\mathcal{C}=\{C_{x_{1},\dots,x_{n}}:x_{1},\dots,x_{n}\in\mathscr{X},n>1\}$. Since $\sigma$ is the c.d.f. of a standard logistic random variable, the collection of copula functions satisfies conditions of Corollary 1 of Barrientos et al., (2012). Also, the stick-breaking ratios of logistic-beta dependent Dirichlet process $\sigma\{\eta_{h}(x)\}$ are determined by the set of copulas $\mathcal{C}$ and the set of beta marginal distributions with parameters $1$ and $b$. Therefore, the full weak support property of the logistic-beta dependent Dirichlet process follows from Theorem 3 of Barrientos et al., (2012). For a general logistic-beta dependent Dirichlet process where atoms and weights both depend on $x$, the full weak support condition is satisfied as long as the atom process can be represented with a collection of copulas with positive density w.r.t. Lebesgue measure.
    S.1.3.2 Proof of Theorem 4.2 and Theorem 4.3